# Boosting CLIP Adaptation for Image Quality Assessment via Meta-Prompt Learning and Gradient Regularization

### v1

@@ Abstract
[…]
@@ 1 Introduction
Figure 1: Intuitive diagram of GRMP-IQA. (a) demonstrates using meta-learning for efficient soft prompt initialization with quality prior that enhances zero-shot generalization ($\uparrow 24\%$), enabling the CLIP to adapt to BIQA tasks effectively. (b) shows a gradient regularization that fine-tunes the model by regularizing between quality and semantic gradients for IQA tasks, reducing the model’s overemphasis on semantic features. The black dots in t-SNE visualization represent noise samples with high semantic confidence ($\text{confidence}\bm{>0.8}$) but incorrect quality predictions. The noticeable decrease ($\downarrow 73\%$) in black dots indicates a reduction in semantic noise interference with quality assessments.
@@
With the onset of the mobile internet era, the focus on computer vision has transitioned from initial concerns with compression and image processing (Sheikh, Sabir, and Bovik 2006) to handling user-generated content like smartphone photos and videos (Tu et al. 2021), and lately to AI-generated content (Li et al. 2023a). This progression has precipitated a marked escalation in the need for effective Blind Image Quality Assessment (BIQA) techniques, underscoring the importance of developing methodologies that can adeptly evaluate image quality without reference images.
Data-driven BIQA models (Song et al. 2023; Yang et al. 2022) based on deep neural networks have made significant progress in recent years. However, the quality scores for distorted images are often measured using the Mean Opinion Score (MOS), which is the average of multiple ratings (sometimes up to 120). As a result, acquiring a sufficient number of IQA training samples is quite labor-intensive and cost-expensive. Recent BIQA approaches have tackled this challenge by leveraging the rich representations of large-scale datasets. Studies like (Qin et al. 2023; Xu et al. 2024) employ transfer learning by pretraining a model on ImageNet (Deng et al. 2009). Some works (Saha, Mishra, and Bovik 2023; Zhao et al. 2023; Srinath et al. 2024) propose IQA-specific pretext tasks, assuming that image distortions directly affect quality. These methods aim to extract quality representations efficiently through self-supervised learning, using scalable datasets. However, they often require extensive pre-training and involve large-scale networks with millions of parameters, leading to high computational costs.
Recently, the pre-trained vision-language model CLIP (Radford et al. 2021) has achieved notable success in various downstream tasks. CLIP-IQA (Wang, Chan, and Loy 2023) shows promising zero-shot generalization in IQA tasks by using simple handcrafted prompts (e.g., “a good photo”). However, slight changes in these prompts can significantly impact performance, requiring experts to spend time on extensive validation. To overcome this, recent methods (Zhou et al. 2022b; Zhu et al. 2023; Jia et al. 2022) propose learning soft prompts (continuous embeddings) from small amounts of labeled data, offering a promising alternative to manual prompt design.
Despite the clear improvements in downstream task performance, prompt tuning for small labeled data generalization in IQA still faces two major limitations: (1) Sensitivity to Initialization: The performance of soft prompts is highly dependent on their initial settings, especially when training samples are limited. As shown in Fig. 2, there is a significant variation in the average SRCC performance across different initializations, necessitating careful tuning for each new IQA scenario, thereby limiting rapid adaptation. (2) Reduced Generalization: The CLIP encoder is designed for semantic content (e.g., identifying cats or dogs) rather than image quality attributes (e.g., noise or sharpness). When fine-tuned with a small sample size, the model tends to overfit to spurious correlations (such as semantic noise) rather than true image quality. As illustrated in Fig. 1, when the CLIP encoder is fine-tuned on a limited dataset, the latent space contains many instances (black dots) that confidently and accurately predict image semantics (with confidence levels above 0.8) but fail to correctly assess image quality.
@@
To address these limitations, we propose the Gradient-Regulated Meta-Prompt learning for IQA (GRMP-IQA), which operates in two main stages: (1) Meta Prompt Pre-training and (2) Quality-Aware Gradient Regularization. In the first stage, we refine the initialization of soft prompts to reduce their sensitivity and improve adaptability to new IQA scenarios. We design meta-training tasks using a well-annotated dataset with various distortions (e.g., overexposure, blur) and employ a bi-level gradient descent method. This allows the model to generalize across different IQA tasks by learning shared meta-knowledge of quality across distortions. In the second stage, the QGR module modulates gradient updates during fine-tuning by balancing the similarity between quality knowledge direction $G_{\text{qua}}$ and semantic knowledge direction $G_{\text{sem}}$. By clipping the quality gradient $G_{\text{qua}}$ along the semantic gradient $G_{\text{sem}}$, the module reduces reliance on semantic content, thereby enhancing the accuracy of image quality assessments. Moreover, fine-tuning only CLIP’s text branch can misalign quality perception between the image and text branches, limiting generalization (Khattak et al. 2023a). To resolve this, we integrate text prompt tuning (CoOp (Zhou et al. 2022b)) and visual prompt tuning (VPT (Jia et al. 2022)) by jointly meta-learning initializations for both, ensuring complementary optimization to better adapt CLIP to new IQA scenarios (Tab. 4).
@@
Figure 2: The fine-tuning results on the LIVEC dataset using CoOp (Zhou et al. 2022b) demonstrate that the accuracy of prompt tuning is significantly influenced by the initial random initialization of the prompt. This effect becomes more evident with limited training data, emphasizing the importance of prompt initialization under data constraints.
@@
  • Extensive empirical evidence underscores the effectiveness and efficiency of our approach. Remarkably, with only 200 data samples, our method surpasses SOTA models on the LIVEC dataset (using 20% of the training data) and achieves highly competitive performance on the KonIQ dataset (using 2% of the training data).
@@ 2.1 Deep Learning Based BIQA methods
Early CNN-based methods utilized standard pre-training and fine-tuning pipelines (Zhang et al. 2018; He et al. 2016), while meta-learning approaches like MetaIQA (Zhu et al. 2020) improved adaptability from synthetic to real-world images. However, CNNs struggle with non-local features, a gap filled by ViT-based approaches (Golestaneh, Dadsetan, and Kitani 2022; Ke et al. 2021). Recent models like LIQE (Zhang et al. 2023) integrate CLIP’s multi-task learning capabilities, enhancing IQA with supervised fine-tuning across datasets. Yet, these methods heavily rely on extensive annotated data, posing significant cost and time challenges. In contrast, limited-data BIQA models remain underexplored. DEIQT (Qin et al. 2023) shows that fine-tuning a pre-trained ViT model can yield good results with less annotation, though improvement is still needed. Self-supervised methods (Saha, Mishra, and Bovik 2023; Zhao et al. 2023; Srinath et al. 2024) attempt to address data limitations by devising pre-training tasks aligned with BIQA. However, they require large pretext datasets and extensive training, leading to high costs. In contrast, our approach uses prompt tuning to significantly reduce pre-training expenses. By leveraging the synergy between pre-trained CLIP model insights and IQA specifics during fine-tuning, our model is poised to outperform fully supervised BIQA with minimal data.
@@ 2.2 Prompt Tuning
Inspired by Natural Language Processing, recent research has adapted vision-language pre-trained models to downstream tasks by learning task-specific prompts in an end-to-end manner (Wu and Shi 2022; Vu et al. 2021). Given the limited labeled data during training, prompt tuning is often treated as a few-shot learning task (Fu, Fu, and Jiang 2021; Zhang et al. 2024; Li et al. 2023b). CoOp (Zhou et al. 2022b) optimizes prompt vectors in the CLIP language branch for task adaptation, though it struggles with generalization to novel tasks. To address this, CoCoOp (Zhou et al. 2022a) uses a lightweight meta-network to generate input-conditioned tokens, improving adaptability. KgCoOp (Yao, Zhang, and Xu 2023) further enhances generalization by narrowing the gap between handcrafted and learnable prompts. While these methods focus on textual prompts, significant work has also explored visual prompts for task adaptation (Jia et al. 2022). By incorporating trainable prompts into both the language and visual branches of CLIP, methods like MaPLe (Khattak et al. 2023a) and PromptSRC (Khattak et al. 2023b) achieve substantial performance improvements across base and novel tasks.
Figure 3: The overview of our GRMP-IQA. The core modules are the Meta-Prompt Pre-training and Quality-Aware Gradient Regularization, corresponding to two training processes. At the pre-training stage, we establish a distortion Meta-Knowledge task for BIQA tasks, and the bi-level gradient descent is utilized to train visual-text meta-prompt $\bm{[\theta_{T},\theta_{V}]}$ (Sec. 3.3). These optimized prompts then serve as initial settings to efficiently adapt the CLIP model to IQA tasks with limited labels. At the fine-tuning stage, we adjust IQA task gradients $\bm{G_{\text{qua}}}$ by clipping gradients aligned with semantic task gradients $\bm{G_{\text{sem}}}$ to generate $\bm{G_{\text{qgr}}}$ for backward updates (Sec. 3.4), thereby mitigating the influence of semantic noise on quality predictions.
@@ 3.1 Overview
In this paper, we introduce the Gradient-Regulated Meta-Prompt Image Quality Assessment (GRMP-IQA), which aims to adapt the CLIP for BIQA tasks with a few training samples. As depicted in Fig. 3, GRMP-IQA consists of two primary module: the Meta-Prompt Pre-training Module (MPP) and the Quality-Aware Gradient Regularization (QGR). The MPP module pre-trains visual-text prompts to acquire shared meta-knowledge on distortions, enabling quick adaptation to various IQA scenarios. The QGR module plays a key role in fine-tuning by adjusting gradient updates to prevent overfitting to semantic content.
Pre-training stage (Sec. 3.3). We randomly sample a mini-batch from distortion meta-tasks, partitioning it into a support set $\mathcal{D}^{\text{support}}$ and a query set $\mathcal{D}^{\text{query}}$. A bi-level gradient descent method progresses from the Inner-Loop on $\mathcal{D}^{\text{support}}$ to the Outer-Loop on $\mathcal{D}^{\text{query}}$ to optimize the learnable visual-textual prompts $[\bm{\theta_{T}},\bm{\theta_{V}}]$. These prompts are then used as initial conditions for fine-tuning.
Fine-tuning stage (Sec. 3.4). Given an input image $x$ and a semantic prompt $w_{i}$ =“a photo of a [class].”, we predict the class probability distribution $\bm{p^{sem}(w_{i}|x)}$ and $\bm{p^{qua}(w_{i}|x)}$ of image $x$ using both the origin semantic CLIP model and fine-tuned CLIP model and compute the gradient of the KL divergence loss $\nabla\mathcal{L}_{kl}$ as a measure of general semantic direction. Simultaneously, according to prompt $\bm{\theta_{T}}$ and image $x$, we compute the gradient $\nabla\mathcal{L}_{ce}$ of the quality loss between the predicted quality and the actual quality label as the quality direction. Finally, we refine the IQA task direction $\bm{G}_{\text{qua}}\rightarrow\bm{G}_{\text{qgr}}$ by clipping gradients aligned with the general semantic direction $\bm{G}_{\text{sem}}$. This approach ensures that the model prioritizes image quality, thereby minimizing the impact of semantic noise.
@@ 3.2 Visual-Text Meta-Prompt
Visual Meta-Prompt. We use Deep Prompt Tuning (DPT (Zhou et al. 2022b)) as our Visual Meta-Prompt. The input embedding for the $\bm{l}$ layer’s self-attention module in the ViT-based image encoder is denoted as $\bm{\{f^{\bm{l}},h_{1}^{\bm{l}},h_{2}^{l},\ldots,h_{N}^{l}}\}$, where $\bm{f^{\bm{l}}}$ represents the classification (CLS) token, and $\bm{H^{\bm{l}}}$ denotes the image patch embeddings. A learnable token $\bm{P^{\bm{l}}}$ is appended to the token sequence in each ViT layer, and the Multi-Head Attention module (MHA) processes the tokens as:
$[{f^{l}},\_,{H^{l}}]=\text{Layer}^{l}([{f^{{l-1}}},{P^{l-1}},{H^{l-1}}]),$ (1)
@@ 3.3 Meta-Prompt Pre-training Module
During meta-learning, we optimize the learnable visual-text prompt $\bm{\theta=[{\theta_{T}},{\theta_{V}}]}$. For an input image $\bm{x}$, the textual encoder $\bm{g(\cdot)}$ processes the quality prompts ${q_{i}}$, while the visual encoder extracts the feature vector $\bm{f}$. The probability of predicting high quality is:
$p(q_{\text{high}}|x)=\frac{\exp(\langle g(q_{\text{high}}),f\rangle/\tau)}{\exp(\langle g(q_{\text{high}}),f\rangle/\tau)+\exp(\langle g(q_{\text{low}}),f\rangle/\tau)},$ (2)
where $q_{high}$ and $q_{low}$ donate high quality and low quality category prompt, ${p(q_{\text{high}}|x)}$ represents the estimated probability that the quality of image $\bm{x}$ is “high quality”, $\bm{\tau}$ is a temperature parameter learned by CLIP, and $\bm{\left\langle,\right\rangle}$ denotes cosine similarity. The labeled quality scores are then rescaled to 0-1, denoted as $y$, and the loss function is calculated as:
$\mathcal{L}=-(y\log(p(q_{\text{high}}|x))+(1-y)\log(1-p(q_{\text{high}}|x))).$ (3)
@@
Constructing Distortion Meta-Knowledge Task. As noted by (Ma et al. 2019; Zhu et al. 2020), the ability to detect various types of image distortions is crucial for developing BIQA models with strong generalization capabilities across diverse scenarios. Moreover, the efficacy of prompt tuning is significantly dependent on the initial configuration of the prompts. This initial setup greatly affects the CLIP vision-language model’s ability to swiftly adapt to different IQA scenarios. Drawing inspiration from the “learn to learn” ethos inherent in the deep meta-learning paradigm (Vanschoren 2018; Zhu et al. 2020), we propose an optimization-based method for effectively pre-training visual-text prompts. These prompts incorporate shared quality insights from various image distortions, thereby enhancing CLIP’s swift adaptability to IQA tasks. To investigate the general rules of image distortion, we first constructed a $K_{t}$-way task-specific BIQA task, denoted as $\mathcal{T}_{t}$. Then, it is used to build the meta-training set as $\mathcal{D}^{\text{meta}}=\left\{\mathcal{D}_{\mathcal{T}_{t}}^{\text{support}},\mathcal{D}_{\mathcal{T}_{t}}^{\text{query}}\right\}_{t=1}^{T}$. Here, $\mathcal{D}_{\mathcal{T}_{t}}^{\text{support}}$ and $\mathcal{D}_{\mathcal{T}_{t}}^{\text{query}}$ represent the support and query sets for each task, respectively, with $T$ representing the total number of tasks. To simulate the process of prompt generalization to different distortions in BIQA scenarios, we randomly sample $k$ tasks as a mini-batch from the meta-training set, where $1\leq k\leq T$, to perform bi-level gradient optimization.
@@
Distortion-Aware Meta-prompt Learning. Our approach employs a bi-level gradient descent technique to bridge the learning process from the support to the query set. Specifically, it mainly consists of two optimization steps. In the Inner-Loop (the first level), we compute the gradients of the prompt parameters using the support set and apply the first update. In the Outer-Loop (the second level), we assess the performance of the updated model on the query set. This bi-level structure enables the prompt to rapidly acquire generalization ability across diverse BIQA contexts by optimizing meta-prompts to acquire shared quality prior knowledge.
@@
Inner-Loop. The objective of the Inner-Loop stage is to adapt the meta-prompt, denoted by $\bm{\theta}=[{\bm{\theta}_{T}},{\bm{\theta}_{V}}]$, to the $t^{th}$ support set ${\mathcal{D}}_{{\mathcal{T}t}}^{\text{support}}$ within the mini-batch. During the first level of updates, we determine the loss $\mathcal{L}(\theta,\mathcal{D}_{\mathcal{T}_{t}}^{\text{support}})$, following which the model parameters are updated on the support set using the inner learning rate $\alpha$, as specified by:
$\theta^{\prime}_{i}\leftarrow\theta-\alpha{\nabla\theta\mathcal{L}(\theta,\mathcal{D}_{\mathcal{T}_{t}}^{\text{support}})},$ (4)
@@
Outer-Loop. In a similar vein, the second level of updates adjusts the parameters $\bm{\theta^{\prime}_{i}}$ based on the query set ${\mathcal{D}_{\mathcal{T}_{t}}^{\text{query}}}$:
@@
$\theta_{i}\leftarrow\theta_{i}^{\prime}-\alpha{\nabla\theta\mathcal{L}(\theta,\mathcal{D}_{\mathcal{T}_{t}}^{\text{query}})}.$ (5)
@@
The culmination of this process for a mini-batch of meta-tasks involves the aggregation of gradients from all tasks to update the final model parameters, following the update rule:
$\theta\leftarrow\theta-\beta\frac{1}{k}\sum\limits_{i=1}^{k}{(\theta-{\theta_{i}})},$ (6)

### v2

# Few-Shot Image Quality Assessment via Adaptation of Vision-Language Models

@@ Abstract
Image Quality Assessment (IQA) remains an unresolved challenge in computer vision due to complex distortions, diverse image content, and limited data availability. Existing Blind IQA (BIQA) methods largely rely on extensive human annotations, which are labor-intensive and costly due to the demanding nature of creating IQA datasets. To reduce this dependency, we propose the Gradient-Regulated Meta-Prompt IQA Framework (GRMP-IQA), designed to efficiently adapt the visual-language pre-trained model, CLIP, to IQA tasks, achieving high accuracy even with limited data. GRMP-IQA consists of two core modules: (i) Meta-Prompt Pre-training Module and (ii) Quality-Aware Gradient Regularization. The Meta Prompt Pre-training Module leverages a meta-learning paradigm to pre-train soft prompts with shared meta-knowledge across different distortions, enabling rapid adaptation to various IQA tasks. On the other hand, the Quality-Aware Gradient Regularization is designed to adjust the update gradients during fine-tuning, focusing the model’s attention on quality-relevant features and preventing overfitting to semantic information. Extensive experiments on standard BIQA datasets demonstrate the superior performance to the state-of-the-art BIQA methods under limited data setting. Notably, utilizing just 20% of the training data, GRMP-IQA is competitive with most existing fully supervised BIQA approaches. Our code is available via https://github.com/LXDxmu/GRMP-IQA.
*Equal contribution.  Corresponding authors.
@@ 1 Introduction
Figure 1: Intuitive diagram of GRMP-IQA. (a) demonstrates using meta-learning for efficient soft prompt initialization with quality prior that enhances zero-shot generalization ($\uparrow 24\%$), enabling the CLIP to adapt to BIQA tasks effectively. (b) illustrates gradient regularization during fine-tuning, which clips quality gradients aligned with semantic bias to guide the model toward quality-relevant features and reduce overly reliance on semantic content. In the t-SNE visualization, black dots represent noise samples with high semantic confidence ($\text{confidence}\bm{>0.8}$) but incorrect quality predictions. The noticeable decrease ($\downarrow 73\%$) in black dots highlights reduced semantic noise interference in IQA tasks.
@@
With the rise of the mobile internet era, the focus on computer vision has transitioned from initial concerns with compression and image processing [38] to handling user-generated content like smartphone photos and videos [42, 6, 35], and lately to AI-generated content [20, 5]. This evolution has sharply increased demand for effective Blind Image Quality Assessment (BIQA) techniques, highlighting the importance of developing methodologies that can adeptly evaluate image quality without reference images. Data-driven BIQA models [39, 48, 34] based on deep neural networks have made significant progress in recent years. However, the quality scores for distorted images are often measured using the Mean Opinion Score (MOS), which is the average of multiple ratings (sometimes up to 120). As a result, acquiring a sufficient number of IQA training samples is quite labor-intensive and cost-expensive.
To mitigate this challenge, recent BIQA approaches [33, 46, 37, 53, 40] have leveraged large-scale image datasets through transfer learning or self-supervised learning. While these approaches improve performance, they often rely on computationally intensive pretraining or require domain-specific pretext tasks, limiting scalability. On the other hand, the vision-language model CLIP [36] has shown remarkable generalization across various downstream tasks, offering a potential alternative. Specifically, CLIP-IQA [44] demonstrates promising zero-shot performance in IQA by leveraging simple handcrafted prompts (e.g., “Good/Bad photo”). However, a significant performance gap remains compared to fully supervised methods [40, 45], highlighting the need to leverage limited labeled data to further close this performance gap. To bridge this gap, we introduce the concept of Few-Shot IQA, which aims to enable models to quickly generalize to new distortion types and unseen scenarios with only a small amount of labeled data (e.g., 50–200 training images). A natural way to adapt CLIP to downstream IQA tasks under this setup is through prompt tuning [56, 57, 11]. Prompt tuning techniques learn soft textual prompts (continuous embeddings) from a small labeled dataset while keeping the pretrained model parameters frozen, offering a lightweight and efficient solution.
However, as shown in Tab. 1, the direct application of the prompt tuning method CoOp [56] to IQA yields minimal improvement over simple linear probing on the CLIP image encoder, falling short of the expected benefits seen in high-level vision tasks. We attribute this limitation to a fundamental mismatch between CLIP’s pretraining objectives and the specific requirements of IQA tasks. CLIP is pretrained to align text and images in a shared semantic space, with a focus on high-level semantics such as objects and scenes. In contrast, IQA demands fine-grained sensitivity to low-level distortion cues—an ability in which CLIP is relatively deficient [52, 27]. This discrepancy introduces two key challenges when applying prompt tuning to IQA tasks:
(1) Sensitivity to Initialization: During prompt tuning, with CLIP’s parameters frozen and limited sensitivity to distortions, the model depends heavily on the learned prompts for quality assessment. Different prompt initializations lead to distinct optimization directions [18], which may steer optimization toward features less relevant to quality, causing significant performance variability. As shown in Fig. 2, the average SRCC fluctuates considerably across different random initializations, requiring careful tuning for each IQA scenario and limiting the model’s adaptability.
(2) Overfitting from Semantic Bias: Due to CLIP’s inherent reliance on semantic information, fine-tuning with a small sample size often leads to overfitting, where the model learns spurious correlations (e.g., over-reliance on semantic features to infer quality [22, 1]) rather than true image quality indicators, which degrades IQA performance. As illustrated in Fig. 1, when the CLIP encoder is fine-tuned on a limited dataset, the latent space contains many instances (black dots) that confidently and accurately predict image semantics (with confidence levels above 0.8) but fail to reliably assess image quality.
@@
To overcome these limitations, we propose Gradient-Regulated Meta-Prompt Learning for IQA (GRMP-IQA), a framework with two key components: (i) Meta Prompt Pre-training (MPP) mitigates sensitivity to prompt initialization by incorporating generalized distortion priors into the initialization of soft prompts, which enhances their adaptability to new IQA scenarios. By designing meta-training tasks with a well-annotated dataset containing diverse distortions (e.g., overexposure, blur) and optimizing prompts through bi-level gradient descent, we enable the model to learn shared meta-knowledge of quality representations across different distortions, thereby improving generalization across IQA tasks. (ii) Quality-Aware Gradient Regularization (QGR) addresses overfitting to semantic information by regulating gradient updates during fine-tuning, which balances the interplay between the quality knowledge gradient $\bm{G}_{\text{qua}}$ and the semantic knowledge gradient $\bm{G}_{\text{sem}}$. By clipping $\bm{G}_{\text{qua}}$ along $\bm{G}_{\text{sem}}$ when their directions are overly aligned, this module suppresses semantic bias and guides the model to focus on distortion-related features, improving its ability to capture low-level quality details. Additionally, fine-tuning only CLIP’s text branch can misalign quality perception between the image and text branches, limiting generalization [14]. To address this, we integrate text prompt tuning (CoOp [56]) and visual prompt tuning (VPT [11]) by jointly meta-learning initializations for both, ensuring complementary optimization to better adapt CLIP to new IQA scenarios (Tab. 5).
@@
Figure 2: The fine-tuning results on the LIVEC dataset using CoOp [56] show that prompt tuning accuracy is highly sensitive to initial random initialization of the prompt, especially with limited training data, underscoring its critical role under data constraints.
@@
  • Extensive empirical results confirm that our approach is both effective and efficient. Notably, with just 200 data samples, our method is competitive with SOTA models on the LIVEC dataset (using 20% of the training data).
@@ 2.1 Deep Learning Based BIQA methods
Early CNN-based IQA methods followed standard pre-training and fine-tuning pipelines [51, 9], while meta-learning approaches like MetaIQA [58] improved adaptability from synthetic to real-world images. However, CNNs struggle with non-local features, a gap addressed by ViT-based models [8, 13, 23, 24]. Recent methods like LIQE [52] leverage CLIP’s multi-task learning, enhancing IQA through supervised fine-tuning across datasets. While effective, these methods rely heavily on extensive annotations, making them costly and time-intensive. In contrast, limited-data BIQA remains underexplored. CLIP-IQA[44] shows promise in zero-shot settings but has performance limits, highlighting the need for better use of limited labeled data. Methods like DEIQT[33] and LoDa[46] demonstrate that fine-tuning pre-trained ViTs reduces annotation needs but still underperform in few-shot scenarios. Self-supervised approaches[37, 53, 40] ease data constraints through tailored pre-training tasks but carry high computational costs. Additionally, recent studies [22, 1] explore disentangling semantic and distortion-related content for IQA—yet none fully address the dual challenges of data efficiency and computational cost.
To bridge these gaps, our approach combines CLIP’s pre-trained semantic knowledge with distortion-specific fine-tuning. By using prompt tuning to cut pre-training costs and a parameter-free regularization strategy (leveraging high- and low-level gradient correlations), it optimizes IQA more effectively with fewer labeled samples.
@@ 2.2 Prompt Tuning
Given the limited labeled data during training, prompt tuning is an effective approach to adapt vision-language pre-trained models for few-shot learning tasks [4, 50, 47]. CoOp [56] optimizes prompt vectors in CLIP’s language branch for task adaptation but struggles with generalization to novel tasks. To address this, CoCoOp [55] introduces a lightweight meta-network to generate input-conditioned tokens, improving adaptability. Beyond textual prompts, visual prompts have also been explored for task adaptation [11]. Methods like MaPLe [14] and PromptSRC [15] incorporate trainable prompts into both language and visual branches, achieving significant performance gains across base and novel tasks. Recent studies [31, 21] further construct meta-tasks based on semantic categories to initialize well-structured soft prompts, mitigating generalization decline in high-level tasks. Unlike these approaches that focus on high-level tasks, our proposed prompt tuning method is specifically designed for IQA, leveraging low-level distortion information to construct meta-tasks. This enables pre-trained prompts to be distortion-sensitive and effectively capture shared quality knowledge.
Figure 3: The overview of our GRMP-IQA. The core modules are the Meta-Prompt Pre-training and Quality-Aware Gradient Regularization, corresponding to two training processes. At the pre-training stage, we establish a distortion-specific meta-knowledge task for BIQA tasks, and the bi-level gradient descent is utilized to train visual-text meta-prompt $\textbf{[}\bm{\theta_{T}},\bm{\theta_{V}}\textbf{]}$ (Sec. 3.3). These optimized prompts then serve as initial settings to efficiently adapt the CLIP model to IQA tasks with limited labels. At the few-shot fine-tuning stage, we first predict the class probability distributions $\bm{p^{sem}(w_{i}|x)}$ and $\bm{p^{qua}(w_{i}|x)}$ using the semantic and fine-tuned CLIP models, respectively. We then adjust IQA task gradients $\bm{G_{\text{qua}}}$ by clipping gradients aligned with semantic task gradients $\bm{G_{\text{sem}}}$ to generate refined gradient $\bm{G_{\text{qgr}}}$ for backward updates (Sec. 3.4), mitigating the influence of semantic noise on quality predictions.
@@ 3.1 Overview
In this paper, we propose the Gradient-Regulated Meta-Prompt Image Quality Assessment (GRMP-IQA), which aims to adapt the CLIP for BIQA tasks with a few training samples. As depicted in Fig. 3, GRMP-IQA consists of two primary module: (i) the Meta-Prompt Pre-training Module (MPP) and (ii) the Quality-Aware Gradient Regularization (QGR). The MPP module pre-trains visual-text prompts to acquire shared meta-knowledge on distortions, enabling quick adaptation to various IQA scenarios. The QGR module plays a key role in fine-tuning by adjusting gradient updates to prevent overfitting to semantic content.
Pre-training stage (Sec. 3.3). We randomly sample a mini-batch from distortion meta-tasks, partitioning it into a support set $\mathcal{D}^{\text{support}}$ and a query set $\mathcal{D}^{\text{query}}$. A bi-level gradient descent method progresses from the Inner-Loop on $\mathcal{D}^{\text{support}}$ to the Outer-Loop on $\mathcal{D}^{\text{query}}$ to optimize the learnable visual-textual prompts $\textbf{[}\bm{\theta_{T}},\bm{\theta_{V}}\textbf{]}$. These prompts are then used as initial weights for following fine-tuning.
[…]
@@ 3.2 Visual-Text Meta-Prompt
Visual Meta-Prompt. We utilize Deep Prompt Tuning (DPT) [56] as our Visual Meta-Prompt, with learnable parameters $\bm{\theta_{V}}$. The input embedding for the $l$-th layer’s self-attention module in the ViT-based image encoder is denoted as $\{\bm{f}^{l},\bm{H}^{l}\}$, where $\bm{f}^{l}$ represents the classification (CLS) token, and $\bm{H}^{l}=\{\bm{h}_{1}^{l},\bm{h}_{2}^{l},\ldots,\bm{h}_{N}^{l}\}$ denotes the image patch embeddings. A learnable token $\bm{P}^{l}$ is appended to the token sequence in each ViT layer, and the Multi-Head Self-Attention (MHSA) module processes the tokens as:
$\bm{f}^{l},\_,\bm{H}^{l}]=\text{Layer}^{l}([\bm{f}^{{l-1}},\bm{P}^{l-1},\bm{H}^{l-1}]),$ (1)
@@ 3.3 Meta-Prompt Pre-training Module
During meta-learning, we aim to optimize the learnable visual-text prompt $\bm{\theta=[{\theta_{T}},{\theta_{V}}]}$. For an input image $\bm{x}$, the textual encoder $\bm{g(\cdot)}$ processes the quality prompts $\bm{q}_{i}$, while the visual encoder extracts the feature vector $\bm{f}$. The probability of predicting high quality is:
$p(\bm{q}_{\text{high}}|\bm{x})=\frac{\exp(\langle\bm{g}(\bm{q}_{\text{high}}),\bm{f}\rangle/\bm{\tau})}{\exp(\langle\bm{g}(\bm{q}_{\text{high}}),\bm{f}\rangle/\bm{\tau})+\exp(\langle\bm{g}(\bm{q}_{\text{low}}),\bm{f}\rangle/\bm{\tau})},$ (2)
[…]
$\mathcal{L}=-(y\log(p(q_{\text{high}}|\bm{x}))+(1-y)\log(1-p(q_{\text{high}}|\bm{x}))).$ (3)
@@
Constructing Distortion Meta-Knowledge Task. As noted by [29, 58], the ability to detect various types of image distortions is crucial for developing BIQA models with strong generalization capabilities across diverse scenarios. Moreover, the efficacy of prompt tuning is significantly dependent on the initial configuration of the prompts. This initial setup greatly affects the CLIP vision-language model’s ability to swiftly adapt to different IQA scenarios. Drawing inspiration from the “learn to learn” ethos inherent in the deep meta-learning paradigm [43, 58], we propose an optimization-based method for effectively pre-training visual-textual prompts. These prompts incorporate shared quality insights from various image distortions, thereby enhancing CLIP’s swift adaptability to IQA tasks [58]. To investigate the general rules of image distortion, we first constructed a $K_{t}$-way distortion-specific image quality prediction task, denoted as $\mathcal{T}_{t}$. Then, it is used to build the meta-training set as $\mathcal{D}^{\text{meta}}=\left\{\mathcal{D}_{\mathcal{T}_{t}}^{\text{support}},\mathcal{D}_{\mathcal{T}_{t}}^{\text{query}}\right\}_{t=1}^{T}$. Here, $\mathcal{D}_{\mathcal{T}_{t}}^{\text{support}}$ and $\mathcal{D}_{\mathcal{T}_{t}}^{\text{query}}$ represent the support and query sets for each task, respectively, with $T$ representing the total number of tasks. To simulate the process of prompt generalization to different distortions in BIQA scenarios, we randomly sample $k$ tasks as a mini-batch from the meta-training set, where $1\leq k\leq T$, to perform bi-level gradient optimization.
@@
Distortion-Aware Meta-prompt Learning. Our approach employs a bi-level gradient descent technique to bridge the learning process from the support to the query set. Specifically, it mainly consists of two optimization steps. In the Inner-Loop (the first level), we compute the gradients of the prompt parameters using the support set and apply the first update. In the Outer-Loop (the second level), we assess the performance of the updated model on the query set and optimize the parameters again. This bi-level structure distills shared quality priors by training on a number of NR-IQA tasks with known distortions, enabling the meta-prompt to generalize rapidly across diverse new BIQA scenarios.
@@
Inner-Loop. The objective of the Inner-Loop stage is to adapt the meta-prompt, denoted by $\textbf{[}\bm{\theta_{T}},\bm{\theta_{V}}\textbf{]}$, to the $t^{th}$ support set ${\mathcal{D}}_{{\mathcal{T}_{t}}}^{\text{support}}$ within the mini-batch. During the first level of updates, we determine the loss $\mathcal{L}(\bm{\theta},\mathcal{D}_{\mathcal{T}_{t}}^{\text{support}})$, following which the model parameters are updated on the support set using the inner learning rate $\alpha$, as specified by:
$\bm{\theta}^{\prime}_{t}=\bm{\theta}-\alpha\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta},\mathcal{D}_{\mathcal{T}_{t}}^{\text{support}}),$ (4)
@@
Outer-Loop. In a similar vein, the second level of updates adjusts the parameters $\bm{\theta}^{\prime}_{t}$ based on the query set $\mathcal{D}_{\mathcal{T}_{t}}^{\text{query}}$}:
@@
$\bm{\theta}_{t}\leftarrow\bm{\theta}^{\prime}_{t}-\alpha\nabla_{\bm{\theta}^{\prime}_{t}}\mathcal{L}(\bm{\theta}^{\prime}_{t},\mathcal{D}_{\mathcal{T}_{t}}^{\text{query}}).$ (5)
@@
For a mini-batch of meta-tasks, this process culminates in aggregating gradients from all tasks to update the final model parameters, following the update rule:
$\bm{\theta}\leftarrow\bm{\theta}-\beta\sum_{t=1}^{k}\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta}^{\prime}_{t},\mathcal{D}_{\mathcal{T}_{t}}^{\text{query}}),$ (6)